OpenStudy (anonymous):

Is there a x-intercept? y - 2 = 2(x+1)^2

OpenStudy (anonymous):

Here is a graph, it doesnt look like there is one

OpenStudy (saifoo.khan):

There will be imaginary roots.

OpenStudy (anonymous):

Let's check is there any x for which is y = 0 $$ y = 2(x+1)^2 + 2 = 2((x + 1)^2 + 1) $$ $$ y = 0 \rightarrow (x+1)^2 = -1 $$ which is impossible in set of real numbers

OpenStudy (anonymous):

so it becomes an imaginary number? @fantastic001

OpenStudy (anonymous):

I need to be able to find the: Vertex: (-1, 2) x-intercept: not sure y-intercept: 4 axis o symmetry: -1

OpenStudy (anonymous):

@feegee if x is in set of comples numbers, then it is complex number, in real numbers, x such that y = 0 does not exist because in real numbers square is always positive

OpenStudy (anonymous):

@fantastic001 okay, so from everything I've gotten from you so far- well my understanding, is that there is no x-axis, it doesnt exist.

OpenStudy (anonymous):

I'm sorry, I meant x-intercept**

OpenStudy (anonymous):

yeah, it doesn't exist x-axis intersection because y is never 0 :)

OpenStudy (anonymous):

alright! Do you mind if I bring up another question which is a bit trickier than? @fantastic001

OpenStudy (anonymous):

i'll try to help if i can :)

OpenStudy (anonymous):

@fantastic001 great! It's the same thing, where I need to find the: vertex: x-intercept: y-intercept: -6 axis of symmetry: 1

OpenStudy (anonymous):

okay, so looks like the vertex is : (1, -6.5)

OpenStudy (anonymous):

ok, this is a parabola and equation for parabola is $$ y = ax^2 + bx + c $$ in y-intersection $$ x = 0 $$ so we have $$ y_{intersect} = a*0^2 + b*0 + c = c $$ and you say that $$ y_{intersect} = -6 $$ so $$ c = -6 $$ ok we have $$ c = -6 $$ , we need a and b right? look at graph you have point (1, -6.5) , put that in equation $$ -6.5 = a*1^2 + b -6 $$ $$ 0.5 = a +b $$ OK, , let's put another information, look at graph again, you have point (2, -6) so we have $$ -6 = 4a + 2b - 6 $$ $$ 0 = 4a + 2b $$ $$ 0.5 = a+b $$ it is system of linear equations, we know how to solve it :) $$ b = -2a $$ $$ a = -0.5 $$ $$ b = 1 $$ now make a complete equation $$ y = -0.5*x^2 + x - 6 $$ now solve $$ y = 0 $$ and you'll get intersections with x-axis (there are two intersections !)

OpenStudy (anonymous):

okay, correct me if I'm wrong but I'm getting; \[1 + i \sqrt{11} or 1 - i \sqrt{11}\]

OpenStudy (anonymous):

@fantastic001

OpenStudy (anonymous):

i think you're right

OpenStudy (anonymous):

hmm but hraph on a picture and graph of $$ y = -0.5x^2 + x - 6 $$ are not same hmm

OpenStudy (anonymous):

The function for the graph is in vertex form: y - 2 = 2 (x+1)^2

OpenStudy (anonymous):

@fantastic001

OpenStudy (anonymous):

ok, to find y-intersect just solve $$ y = 2(x + 1)^2 + 2 $$ at $$ x = 0 $$ so $$ y = 4 $$ but this equation has no solution for $$ y = 0 $$ in set of real numbers

OpenStudy (anonymous):

looks like neither of my problems have a solution for the x-intercept.

OpenStudy (anonymous):

for $$ y = 0 $$ you have $$ x = -1 \pm i $$ but this is complex number

OpenStudy (anonymous):

yeah, looks like they both are complex numbers

OpenStudy (anonymous):

for both x-intercepts